OpenStudy (anonymous):

Find the median. 2, 20, 19, 15, 11, 9, 7, 1

OpenStudy (anonymous):

(15+11)/2= 26/2=13

OpenStudy (shinalcantara):

In finding the median all you have to do is find the innermost part of the given data. say if you have the given data as (1,2,3,4,5). by mere inspection, you will find out that 3 is the middle term therefore you'll have the median as 3.. say you have the given as (1,2,3,4).. since the middle terms are 2 and 3, all you have to do is find the average so you'll have (2+3)/2 = 5/2 or 2.5.. now try to examine the given data and find out what's the median :)